OpenStudy (anonymous):

Add and simplify: (4y/y+2) - (8/y+2)

mathslover (mathslover):

\[\huge{\frac{4y}{y+2}-\frac{8}{y+2}}\] \[\huge{\frac{4y-8}{y+2}}\] Remember , whenever the denominators are same and u r said to add / subtract then just put the numerators above only one denominator .. as like this happened here . denominator y+2 is common .. what we did is just put 4y and 8 and subtracted them \[\huge{\frac{4(y-2)}{y-2}}\] \[\huge{4\frac{\cancel{y-2}}{\cancel{y-2}}}\] \[\huge{4}\]

OpenStudy (anonymous):

sorry i meant add and simplify

mathslover (mathslover):

no problem the answer is same . i did mistake there . sorry